OpenStudy (anonymous):

what is 18 divided by 2/3

OpenStudy (anonymous):

@RedGuy46

OpenStudy (anonymous):

sup

OpenStudy (sleepyjess):

Same concept as the last one. \(\sf\dfrac{18}{1}/\dfrac{2}{3}\). Flip the second fraction and multiply ^_^

OpenStudy (anonymous):

okay @sleepyjess

OpenStudy (anonymous):

18*3= 54

OpenStudy (anonymous):

2*1=2

OpenStudy (anonymous):

54/2

OpenStudy (anonymous):

no its not diagonal

OpenStudy (sleepyjess):

Yep, so now you have \(\sf\dfrac{54}{2}\). Then simplify that

OpenStudy (anonymous):

54 divided by 2 = 27

OpenStudy (anonymous):

27/1

OpenStudy (anonymous):

not a option

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

A. 1/12 B. 3 C:12 D: 27

OpenStudy (sleepyjess):

What is 27/1 equal to?